OpenStudy (anonymous):

find two numbers whose product is 64 , and whose sum is a minimim

hartnn (hartnn):

calculus question, right ? do u know how to take derivatives ?

OpenStudy (anonymous):

yes, but i don know how to form an equation

OpenStudy (anonymous):

Ab = 64 Their sum is a minimum. So, the sum S of a and b, S = a + b should be minimum. To get the minimum, you have to find the derivative of S (or find the shape of the graph of S as a and b vary. S = a + b Substitute either a = 64/b or b = 64/a. I'll do the second. S = a + 64/a Now, you can find the derivative of that equation, equate to zero and solve for a, the first number. Then use the first equation to get the value of b.

hartnn (hartnn):

let 2 numbers be x andy xy=64 sum = f(x) = x+y = x+64/x now minimise f(x)

OpenStudy (anonymous):

am i going to derive x + 64/x?

hartnn (hartnn):

yup. cun u ?

hartnn (hartnn):

*can

OpenStudy (anonymous):

1 + 64?

hartnn (hartnn):

nopes, what is the derivative of 1/x ??

OpenStudy (anonymous):

sorry, 1 - 64/x2

hartnn (hartnn):

now equate that to 0

OpenStudy (anonymous):

(x^2 - 64)/x^2 = 0

hartnn (hartnn):

so numerator = 0

OpenStudy (anonymous):

why?

hartnn (hartnn):

if a/b=0 then a=0. (multiplying b on both sides)

hartnn (hartnn):

so here u get x^2-64=0

OpenStudy (anonymous):

yew, then i will transpose -64 to the right side?

hartnn (hartnn):

yup. then take square root.

hartnn (hartnn):

for your reference. x^2=64 x=8 or x=-8 x cannot be 8 because the 2nd derivative becomes negative and u get a maxima. x=-8 is correct because 2nd derivative is positive and u get a minima. so 2 numbers are -8 and 64/(-8) = -8 whose sum = -8-8 = -16